Operations Report – February 20th

Crew 275 Operations Report 20Feb2023
SOL: 8

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet

Notes on non-nominal systems: None

Spirit rover used: Yes
• Hours: 209.1
• Beginning charge: 100%
• Ending charge: 66%
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.1
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 221.0
• Beginning charge: 100%
• Ending charge: 68%
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.5
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report. The EVA Crew forgot a camera lens protector when they came back, on the edge of the trunk of Opportunity rover. Mission Support was advised by radio at the end of the EVA.

Summary of Hab operations: The Crew conducted routine operations. The Lower Deck heater was ON all day. If the external temperature remains the same as today, we will switch it off as of tomorrow.
• WATER USE: 10.30 gallons
• Water: 355.35 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: The EVA Crew experienced an issue with the radio at North Ridge. After issues of communication between the HabCom and the EVA Crew, everyone in the EVA Crew was on Channel 1, and they were unable to communicate between each other for about one minute. Back on channel 1 (EVA leader) and 2 (the 2 other EVA Crew Members), the EVA Crew was able to communicate again. We have no explanation for this, and we will conduct further tests.

Summary of GreenHab operations: The Crew botanist conducted routine operations and planted two microgreens crops to start his experiment. An environmental sensor was placed in the GreenHab.
• WATER USE: 11.10 gallons
• Heater: On
• Supplemental light: From 10:00 pm to 2:00 am
• Harvest: No

Summary of ScienceDome operations: The Crew Scientist conducted routine operations. An environmental sensor was placed in the ScienceDome.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today. An environmental sensor was placed in the RAM.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: A Crew Member slipped in the stairs, and slightly hurt their elbow. The HSO handled the situation and there is nothing to worry about.

Questions, concerns, and requests to Mission Support: None.

Copyright © The Mars Society. All rights reserved. | Main Site